a quadratic relation has zeros at -1 and 3, and a y-intercept of -18. Determine the equation of the relation in vertex form


1
Expert's answer
2021-06-03T14:54:39-0400

We want quadratic in vertex form. In the vertex form quadratic equation is of type

"y = a(x-h)^2 + k"

given that quadratic has zeros at -1 and 3.

so quadratic equation must have factors "(x+1) and (x-3)."

so our quadratic will be of type"y=p(x+1)(x+3)"

now we want value of p.

given that quadratic relation has y intercept of -18.

so when "u=0, y=-18\\

ie 18=p{0+1}{0-3}\\

=-18=p(-3)\\

p=6\\"

so our quadratic relation is "y=6(x+1)(x-3)"

now

"y=6(x+1)(x-3)\\

y=(6x+6)(x-3)=(6x^2-18x+6x-18)\\

y=6x^2-12x-18\\

y=6(x^2-2x-3)\\

y=6(x^2-2x+1-1-3)\\

y=6(x-1)^1-4)\\

y=6(x-1)^2-24\\"

This is the required vertex form.